OpenStudy (virnios):

The ratio of men to women in a certain factory is 3 to 4. There are 228 men. How many workers are there?

OpenStudy (anonymous):

171

OpenStudy (anonymous):

medal pls

OpenStudy (anonymous):

virnios can i have medal

OpenStudy (anonymous):

thx

OpenStudy (kikuo):

that was me

OpenStudy (texaschic101):

532

OpenStudy (virnios):

thank you

OpenStudy (texaschic101):

3/4 = 228/x 3x = 912 x = 304 so if there is 228 men, there is 302 womem...totalling 532